Why zero ring is artinian as well as noetherian?
1
Expert's answer
2020-04-01T10:15:17-0400

Let "\{I_n\}_{n\in\mathbb N}" be ideals of "\{0\}", such that "I_n\subset I_{n+1}" ("I_n\supset I_{n+1}") for every "n". Since "I_n=I_{n+1}=\{0\}" for every "n\ge 1", we obtain that "\{0\}" is a noetherian (an artinian) ring.